OpenStudy (anonymous):

Solve Equation, 2x - 6 = 2(x - 8) + 10 A. x=3 B. x=8 C. Infinitely many solutions D. No solution

OpenStudy (anonymous):

the answer is a

OpenStudy (the_fizicx99):

2x - 6 = 2(x - 8) + 10 Apply distributive property, 2x - 6 = 2x -16 + 10 2x - 6 = 2x -6 They're the same so this is true <--

OpenStudy (anonymous):

thx